# bibliography.bib

### v3

@@ Abstract
Assume that the interval $I=[0,1)$ is partitioned into finitely many intervals $I_{1},\dots,I_{r}$ and consider a map $T\colon I\to I$ so that $T_{|I_{s}}$ is a translation for each $1\leq s\leq r$. We do not assume that the images of these intervals are disjoint. Such maps are called Interval Translation Maps. Let $\mathrm{ITM}(r)$ be the space of all such transformations, where we fix $r$ but not the intervals $I_{1},\dots,I_{r}$, nor the translations. The set $X(T):=\bigcap_{n\geq 0}T^{n}[0,1)$ can be a finite union of intervals (in which case the map is called of finite type), or is a disjoint union of finitely many intervals and a Cantor set (in which case the map is called of infinite type). In this paper we show that there exists an open and dense subset $\mathcal{S}(r)$ of $\mathrm{ITM}(r)$ consisting of stable maps, i.e., each $T\in\mathcal{S}(r)$
@@ 1 Introduction
One of the main aims in the field of dynamical systems is to determine which dynamical systems are stable, and if possible, to prove that such systems are ‘typical’. We will address this question in a setting where systems are never structurally stable, as explained below. For this reason, we define a map to be stable in this context if the next best thing holds: if -loosely speaking- its non-wandering set depends continuously on the system. Note that this notion does not require that the system and its perturbation are conjugate on their non-wandering sets. This paper is concerned with the space of Interval Translation Maps ($\mathrm{ITM}$’s), defined below. In this setting, the non-wandering can be equal to a finite union of intervals, but it can also be equal to a Cantor set. This class generalizes the well-known class of Interval Exchange Transformations ($\mathrm{IET}$’s). For $\mathrm{IET}$’s, the non-wandering set is always the entire interval and the main result of this paper becomes trivial.
@@
That density of structural stability holds for smooth real 1D systems is proven in [MR2342693]; that it holds for complex holomorphic maps in 1D was proven by McMullen and Sullivan [MR1620850]; it does not hold in higher dimension due to examples by Newhouse (and others). That it holds for $\mathrm{ITM}$’s is proved in this paper. This result implies in particular that stable maps are simple: their dynamics on their non-wandering set corresponds to (a union of) circle rotations.
Whether or not a map in $\mathrm{ITM}$ is stable is equivalent to two conditions: absence of critical connections (ACC) and the matching condition (M). Here ACC means that no iterate of a discontinuity (i.e. a critical point) is mapped to another critical point. In the $\mathrm{IET}$ literature this is often called the Keane condition. The condition M means that the return map to a component of the non-wandering set has at most one discontinuity. This condition has no clear analogue for invertible systems. It rules out that one connected component of the non-wandering set consists of two pieces with independent dynamics which join up for ‘coincidental’ reasons.
@@
So in this sense, the results of this paper are the analogue in the $\mathrm{ITM}$ setting of the density of hyperbolicity in the real one-dimensional dynamics (proved in [MR2342693]). The corresponding question for systems on the complex plane (such as quadratic maps) is the Fatou conjecture and is still wide open. Hence the ACC and M conditions can be considered as the analogues of hyperbolicity in our setting.
The main tool in this paper is a transversality result, namely the linear independence of certain critical itinerary vectors. It is of course easy to determine how changing the intervals and translations changes the local dynamics. The challenge is to determine what happens with the global dynamics when we change some chosen local dynamics. To show that we have the required freedom to control the global dynamics, we need to establish the linear independence of certain vectors associated to the orbits of critical points (critical itinerary vectors). The proof of this linear independence relies on a rather detailed and subtle dynamical analysis.
1.1 Interval Translation Mappings
$T|_{I_{i}}(x)=x+\gamma_{i},$
for $1\leq i\leq r$. Note that we must have that $\gamma_{i}\in[-\beta_{i-1},1-\beta_{i}]$ for all $1\leq i\leq r$, since the image of $T$ is contained in $[0,1)$ by definition. These conditions mean that the parameter space $\mathrm{ITM}(r)$ for $\mathrm{ITM}$’s on $r$ intervals is a convex polytope in $\mathbb{R}^{2r-1}$ and we endow $\mathrm{ITM}(r)$ with the corresponding subspace topology.
In the special case when $T$ is bijective, we get the well-known Interval Exchange Transformations ($\mathrm{IET}$’s). Figures 2 and 2 below show examples of an interval translation mapping and of an interval exchange transformation on $4$ intervals.
Figure 1: $\mathrm{ITM}$ on $4$ intervals
Figure 2: $\mathrm{IET}$ on $4$ intervals
For $n\geq 1$, let $X_{n}(T)=T^{n}(I)$ and note that $X_{n+1}\subset X_{n}$. Let
$X=\bigcap^{\infty}_{n=0}X_{n}.$
It is easy to see that for $r=2$ all $\mathrm{ITM}$’s are of finite type and $X$ is a single interval on which the induced map is a rotation. In general, it is known that $\overline{X}$ is equal to the union of two sets $A_{1}$ and $A_{2}$ (one of which can be empty) so that $A_{1}$ is a union of a finite number of intervals and $A_{2}$ is a topological Cantor set (see Lemma 2.1). It was shown in [MR1796167] that $T$ is of finite type if and only if $X$ is a finite union of intervals, and that if $T$ is of infinite type and topologically transitive then $\overline{X}$ is a Cantor set. In Lemma 2.3 we will show that $\overline{X}$ is equal to the non-wandering set of $T$.
The first explicit example of an infinite type $\mathrm{ITM}$ was given in [MR1356616]. In [MR2013352], a special two-parameter family of maps with a renormalisation operator was considered for $r=3$. This operator is very similar to the one obtained by Rauzy induction for $\mathrm{IET}$’s. It was shown in [MR2013352] that within this space, infinite type $\mathrm{ITM}$’s are those that under successive iterations of this renormalisation operator never enter a particular open subset of the parameter space. Moreover, it was shown that the set of infinite type maps has Lebesgue measure zero in parameter space. This result was generalised to the full parameter space $\mathrm{ITM}(3)$ in [MR3124735], also through renormalisation methods. Unfortunately, for larger numbers of intervals, it is far from clear how to find a suitable renormalisation operator, except in some special cases ([MR2013352], [MR2308208], [MR3124735]). Figures 4 and 4 show the parameter space of the special family from [MR2013352]: the right figure (taken from [MR2013352]) shows the set of infinite type maps, while each open triangle in the left figure [^4] The picture by Björn Winckler. corresponds to a region of equivalent maps, in the sense defined in Subsection 7.1, that are all of finite type. These regions are dynamically related through renormalisation, see [MR2013352] for details.
Figure 3: Finite type maps from the family in [MR2013352]
Figure 4: Infinite type maps from the family in [MR2013352]
1.2 Summary of results
These results and pictures suggest that the finite type maps should be prevalent in the parameter space. Our main result shows that this is true in the topological sense, i.e. that the set of all finite type maps contains an open and dense subset of the full parameter space $\mathrm{ITM}(r)$.
@@ 1.2 Summary of results
It is easy to see that stable maps are of finite type (see Corollary 5.2). Moreover, the number of components of $\overline{X}(T)$ is constant in $\mathcal{U}$. The relationship between the number of components of $\overline{X}(T)$ and the stability of $T\in\mathrm{ITM}(r)$ is quite subtle, see Section 7. For example, it is possible to have that the set $\overline{X}(T)$ can jump up or down as is shown in Examples 5.6 and 7.3. So the map $T\mapsto\overline{X}(T)$ is neither upper nor lower semi-continuous.
Let $\mathcal{S}(r)$ be the set of all stable maps in $\mathrm{ITM}(r)$. For a stable map $T$, denote by $[T]$ the connected component containing $T$ in $\mathcal{S}(r)$. We call $[T]$ the stability component of $T$. When we do not need to emphasize a specific map $T$, we call a connected component of $\mathcal{S}(r)$ a stable region. Note that the minimal number $n$ so that $T^{n}(I)=X(T)$ does not need to be constant within a stable region. It is clear that $[T]$ is the maximal neighbourhood $\mathcal{U}$ for which the conditions of stability hold. It is easy to prove that for $r=2$ the entire parameter space $\mathrm{ITM}(2)$ is equal to the closure of a single stability component. On the other hand, $[T]$ is in general not convex, see Example 7.3. The behaviour near the boundary of $[T]$ is surprisingly complex, as we will see in Section 7.
In the space of $\mathrm{IET}$’s, the definition of stability is vacuous, as the non-wandering set is always equal to $[0,1)$, so all of the conditions for stability are trivially satisfied. On the other hand, by the Characterization of Stability Theorem 5.12, every $\mathrm{IET}$ is not stable when considered as a map in the space of $\mathrm{ITM}$’s.
@@
As a corollary, we get the following theorem which shows the prevalence of finite type maps in the parameter space, in the topological sense:
@@
By our Main Theorem I, the stable maps form a dense subset of $\mathrm{ITM}(r)$. Since the set $\mathcal{S}(r)$ is open by definition and stable maps are of finite type, the result follows. ∎
By the Characterization of Stability Theorem 5.12, we know that the return maps to intervals of $X$ must correspond to circle rotations. The stable maps which are also eventually periodic therefore correspond to rational rotations, in the sense that the return map to any component interval of $X$ is a rational rotation.
In Subsection 7.1, we apply our results to the special family from [MR2013352] to show that the connected components of stability are exactly the coloured triangle regions in Figure 4. Moreover, any point not contained in a closure of a triangle is of infinite type. In Subsection 7.2, we will discuss the full parameter space $\mathrm{ITM}(r)$ of $\mathrm{ITM}$’s on $r$ intervals, with the focus on stable regions. In particular, we will discuss examples 7.3 and 7.8 which show that understanding the topology of the stable regions and their boundaries is quite tricky.
@@
For any eventually periodic map $T$, there exists an arbitrarily small perturbation $\tilde{T}$ of $T$ that is of finite type, satisfies the ACC and Matching conditions and is stable by Theorem B.
@@
In Section 3, we prove that the eventually periodic maps are of finite type and that the set of all eventually periodic maps contains a countable dense subset of $\mathrm{ITM}(r)$.
@@ 1.3 Main idea of the proof
The basic problem we deal with everywhere in the paper is: how to perturb a map $T$ so that some of its dynamical properties change, while some others do not? In analogy to smooth one-dimensional dynamics (e.g. [MR1239171]), the dynamical properties of a map depend on the behaviour of critical points. For us, the discontinuities of the map serve as the critical points. Thus the problem is about controlling critical itineraries: we want some of them to change, and some of them to stay the same up to certain return times. Because our maps are piecewise linear, this can be reformulated in terms of the linear independence of certain vectors defined by the itineraries of critical points (see Section 4 for details). The main difficulty is the following: how does the dynamics of a map $T$ prevent the linear dependence of a set of vectors? We deal with this in the proof of Theorem 4.6.
1.4 History of $\mathrm{ITM}$’s, $\mathrm{IET}$’s and the dynamics of piecewise isometries
The subject of interval exchange transformations (IETs) goes back to the 70’s. Some the earliest publications were by Katok [MR0331438], Keane [MR0357739], [MR0435353] and Veech [MR0516048]. As the literature on IETs is huge we will only refer to the survey papers [MR2219821], [MR2648692]. It is well-known that IET correspond to billiards, see for example [MR0399423] and [MR0644840], [MR0832433, Ch. 5] [MR0889254, Ch.2, §4D].
In 1995, Boshernitzan and Kornfeld [MR1356616] introduced interval translation maps as a natural generalisation of $\mathrm{IET}$’s. The main surprise in their paper was an explicit example of an $\mathrm{ITM}$ with three branches which is of infinite type. They also mentioned that $\mathrm{ITM}$’s correspond to billiards with ‘one-sided mirrors’. Such billiards were considered in more detail by [MR3449199]; see Appendix A. In 1993, Levitt [MR1231840] also considered a two-parameter family of $\mathrm{ITM}$’s, related to the holonomy map of certain 1-dimensional foliations in two dimensions. Bruin and Troubetzkoy [MR2013352] showed that one can use a renormalisation approach to study a special two-parameter family of $\mathrm{ITM}$’s quite systematically. Using this approach they show that within their family the set of parameters corresponding to infinite type maps is uncountable and has Lebesgue measure zero. Volk [MR3124735] generalised this result by showing that within the space of $\mathrm{ITM}$’s with three intervals the set of infinite type maps has Lebesgue measure zero. Schmeling and Troubetzkoy [MR1796167] obtained several results for general $\mathrm{ITM}$’s. In [MR2966738] and [MR2152403], considered a special class of $\mathrm{ITM}$’s on $4$ intervals called double rotations.

### v4

# Density of Stable Interval Translation Maps

@@ Abstract
An Interval Translation Mapping (ITM) is a piece-wise translation $T\colon I\to I$ defined on a finite partition $I_{1},\ldots,I_{r}$ of an interval $I$ into $r\geq 2$ sub-intervals. We do not assume that the images of these intervals are disjoint. These maps naturally generalize classical Interval Exchange Transformations (IETs) by removing the bijectivity assumption. Let $\mathrm{ITM}(r)$ be the space of all interval translation mappings, where we fix $r$ but not the intervals $I_{1},\ldots,I_{r}$, nor the translations on each of them. The set $X(T):=\bigcap_{n\geq 0}T^{n}(I)$ is either a finite union of intervals (and $T$ behaves like an IET on those intervals), in which case the map is called of finite type, or is a disjoint union of finitely many intervals and a Cantor set, in which case the map is called of infinite type. In this paper, for an arbitrary $r\geq 2$, we show that the set of finite type maps contains an open and dense set of $\mathrm{ITM}(r)$. This resolves in positive a topological version of a long-standing conjecture due to Boshernitzan and Kornfeld. More precisely, we show that there exists an open and dense subset $\mathcal{S}(r)$ of $\mathrm{ITM}(r)$ consisting of stable maps such that each $T\in\mathcal{S}(r)$
@@ 1 Introduction
Interval Translation Mappings ($\mathrm{ITM}$s) are orientation-preserving piecewise isometries of an interval. As such, they are the generalization of the well-known Interval Exchange Transformations ($\mathrm{IET}$s) obtained by dropping the bijectivity assumption on the mapping. They were introduced by Boshernitzan and Kornfeld in [MR1356616]. The graphs of an $\mathrm{ITM}$ (left) and $\mathrm{IET}$ (right) on $4$ intervals are displayed in the figures below.
Figure 1: $\mathrm{ITM}$ on $4$ intervals
Figure 2: $\mathrm{IET}$ on $4$ intervals
More precisely, an $\mathrm{ITM}$ on $r$ intervals is a map $T\colon[0,1)\to[0,1)$ defined by two sets of parameters:
  • Discontinuities: $10<\beta_{1}<\beta_{2}<\dots<\beta_{r-1}<$ 1;
  • Translation factors: $\gamma_{1},\gamma_{2},\dots,\gamma_{3}$;
so that $T(x)=x+\gamma_{i}$ for $x\in[\beta_{i-1},\beta_{i})$, where $1\leq i\leq r$. The set of all such parameters forms the parameter space $\mathrm{ITM}(r)$ of all Interval Translation Mappings on $r$ intervals. Naturally, $\mathrm{ITM}(r)$ is a closed convex polytope in $\mathbb{R}^{2r-1}$.
The basic difference between $\mathrm{ITM}$s and $\mathrm{IET}$s is the fact that for a typical $\mathrm{ITM}$ $T$ we have that $T(I)\subsetneq I$, where $I=[0,1)$, which implies that we have the following chain of inclusions: $I\supsetneq T(I)\supseteq T^{2}(I)\supseteq T^{3}(I)\supseteq\dots$. This leads to a natural dichotomy for $\mathrm{ITM}$s: A map $T$ is said to be of finite type if there exists an integer $n$ such that $T^{n}(I)=T^{n+1}(I)$, and is of infinite type if $T^{n}(I)\supsetneq T^{n+1}(I)$ for all $n\geq 0$. The first example of an infinite type map was constructed in [MR1356616] by Boshernitzan and Kornfeld. They conjectured that such behavior should be rare:
Boshernitzan–Kornfeld Conjecture.
For all $r\geq 2$, the set of all infinite type $\mathrm{ITM}$s on $r$ intervals is a measure zero subset of $\mathrm{ITM}(r)$.
This conjecture has been the focus of a large part of research in the field of $\mathrm{ITM}$s, but remains widely open, except in a few cases. Namely, it was first established for $r=3$ and a special $2$-parameter family of $\mathrm{ITM}$s in [MR2013352]. It was shown that this family supports a renormalization scheme analogous to the Rauzy–Veech induction for $\mathrm{IET}$s (see also [artigiani2024renormalization]). The dynamical properties of this renormalization imply that the measure of the set of all infinite type maps must be zero, analogous to how the analysis of the Rauzy–Veech induction shows that almost all minimal $\mathrm{IET}$s are uniquely ergodic (see [MR644018], [MR644019]). A similar method of renormalization was used for $r\leq 4$ and a more complicated family of double rotations in [MR2152403] and [MR2966738] (see also [MR4397159]). The conjecture was established in full for $r=3$ in [MR3124735] by showing that almost every $\mathrm{ITM}$ on three intervals can be renormalized to a double rotation. Unfortunately, there has been little progress on developing a renormalization scheme for $\mathrm{ITM}$s on an arbitrary number of intervals (except in some special cases [MR2308208]).
In this paper, instead of using renormalization, we develop a theory of stability for $\mathrm{ITM}$s (Section 5) and show a transversality result that allows for detailed control over global dynamics under perturbation (Section 4, Theorem 4.6). Using these results, we establish the topological version of Boshernitzan–Kornfeld Conjecture, the main result of this paper:
Topological Prevalence of Finite Type Maps.
For every $r\geq 2$, the set of all finite type $\mathrm{ITM}$s on $r$ intervals contains an open and dense subset of $\mathrm{ITM}(r)$.
We say that a map $T$ corresponds to an irrational circle rotation if the return map to every component interval of its non-wandering set is equivalent to an irrational circle rotation. In particular, such maps are of finite type. The methods used in the proof of Topological Prevalence of Finite Type Maps also suggest a significant refinement of the Boshernitzan–Kornfeld Conjecture:
Irrational Rotations Conjecture.
The set of all $\mathrm{ITM}$s on $r$ intervals that correspond to irrational circle rotations has full measure in $ITM(r)$.
It is widely believed that the Boshernitzan–Kornfeld Conjecture is true. In Section 7 we will show that if it is true, then the stability theory developed in this paper implies the Irrational Rotations Conjecture.
We give the formal introduction and statements of our results in the following subsections.
1.1 Interval Translation Mappings
$T|_{I_{i}}(x)=x+\gamma_{i},$
for $1\leq i\leq r$. Note that we must have that $\gamma_{i}\in[-\beta_{i-1},1-\beta_{i}]$ for all $1\leq i\leq r$, since the image of $T$ is contained in $[0,1)$ by definition. These conditions mean that the parameter space $\mathrm{ITM}(r)$ for $\mathrm{ITM}$s on $r$ intervals is a convex polytope in $\mathbb{R}^{2r-1}$ and we endow $\mathrm{ITM}(r)$ with the corresponding subspace topology. In the special case when $T$ is bijective, we get the well-known Interval Exchange Transformations ($\mathrm{IET}$s).
For $n\geq 1$, let $X_{n}(T):=T^{n}(I)$ and note that $X_{n+1}\subset X_{n}$. Let
$X:=\bigcap^{\infty}_{n=0}X_{n}.$
It is easy to see that for $r=2$ all $\mathrm{ITM}$s are of finite type and $X$ is a single interval on which the induced map is a rotation. In general, it is known that $\overline{X}$ is equal to the union of two sets $A_{1}$ and $A_{2}$ (one of which can be empty) so that $A_{1}$ is a union of a finite number of intervals and $A_{2}$ is a topological Cantor set (see Lemma 2.1). It was shown in [MR1796167] that $T$ is of finite type if and only if $X$ is a finite union of intervals, and that if $T$ is of infinite type and topologically transitive then $\overline{X}$ is a Cantor set. In Lemma 2.3 we will show that $\overline{X}$ is equal to the non-wandering set of $T$.
The first explicit example of an infinite type $\mathrm{ITM}$ was given in [MR1356616]. In [MR2013352] for $r=3$ Bruin and Troubetzkoy introduced a special two-parameter family of maps with a renormalisation operator acting on it; we call this family the Bruin–Troubetzkoy family. It was shown that within this family, infinite type $\mathrm{ITM}$s are those that under successive iterations of this renormalisation operator never enter a particular open subset of the parameter space. Moreover, it was shown that the set of infinite type maps has Lebesgue measure zero in parameter space. Figures 4 and 4 show the parameter space of the special family from [MR2013352]: the right figure (taken from [MR2013352]) shows the set of infinite type maps, while each open triangle in the left figure [^4] The picture by Björn Winckler. corresponds to a region of equivalent maps, in the sense defined in Subsection 8.1, that are all of finite type. These regions are dynamically related through renormalisation, see [MR2013352] for details.
Figure 3: Finite type maps from the family in [MR2013352]
Figure 4: Infinite type maps from the family in [MR2013352]
Linear recurrence and weakly mixing properties for the maps the Bruin–Troubetzkoy family were recently studied in [bruin2023interval]. Estimates on the Hausdorff dimension of the set of all infinite type maps in the family were obtained in [artigiani2024renormalization].
A similar analysis was carried out for the more complicated family of double rotations in [MR2152403], [MR2966738] and [MR4397159]. In [MR3124735], it was shown that almost every $\mathrm{ITM}$ on three intervals is of finite type. In [MR2308208], the same result was shown for special families on an arbitrary number of intervals.
This paper initiates a systematic study of $\mathrm{ITM}$s on $r$ intervals without restrictions on $r$.
1.2 Stability of $\mathrm{ITM}$s
@@
One of the main aims in the field of dynamical systems is to determine which dynamical systems are stable, and if possible, to prove that such systems are ‘typical’. It turns out that $\mathrm{ITM}$s are never structurally stable (see Corollary 1.2). For this reason, we define a map to be stable in this context if the next best thing holds: if -loosely speaking- its non-wandering set depends continuously on the system. This notion of stability is weaker than the usual ones in $1$-dimensional dynamics because it does not require that the system and its perturbation are conjugate on their non-wandering sets (see [MR1239171], [MR1312365]). Note that for $\mathrm{IET}$’s, the non-wandering set is always the entire interval, so the main result of this paper becomes trivial.
That density of structural stability holds for smooth real 1D systems is proven in [MR2342693]; that it holds for complex holomorphic maps in 1D was proven by McMullen and Sullivan [MR1620850]; it does not hold in higher dimension due to examples by Newhouse (and others). That it holds for $\mathrm{ITM}$’s is proved in this paper.
@@
Whether or not a map in $\mathrm{ITM}$ is stable is equivalent to two conditions: Absence of Critical Connections (ACC) and the Matching condition (M) (see Section 5 and Theorem 5.12 for more details). Here ACC means that no iterate of a discontinuity (critical point) is mapped to another discontinuity. In the $\mathrm{IET}$ literature this is often called the Keane condition. The Matching condition means that the return map to a component of the non-wandering set has at most one discontinuity. This condition has no clear analogue for invertible systems. It rules out that one connected component of the non-wandering set consists of two pieces with independent dynamics which join up for ‘coincidental’ reasons. This characterization of stability implies in particular that stable maps are simple: their dynamics on their non-wandering set corresponds to (a union of) circle rotations.
@@
So in this sense, the results for $\mathrm{ITM}$s of this paper are the analogue of the density of hyperbolicity in the real one-dimensional dynamics (proved in [MR2342693]). The corresponding question for systems on the complex plane (such as quadratic maps) is the Fatou conjecture and is still wide open. Hence the ACC and M conditions can be considered as the analogues of hyperbolicity in our setting.
@@
It is easy to see that stable maps are of finite type (see Corollary 5.2). Moreover, the number of components of $\overline{X}(T)$ is constant in $\mathcal{U}$. In general, the map $T\mapsto\overline{X}(T)$ is neither upper nor lower semi-continuous. For example, it is possible to have that the set $\overline{X}(T)$ can jump up or down as is shown in Example 5.6.
Let $\mathcal{S}(r)$ be the set of all stable maps in $\mathrm{ITM}(r)$. For a stable map $T$, denote by $[T]$ the connected component containing $T$ in $\mathcal{S}(r)$. We call $[T]$ the stability component of $T$. When we do not need to emphasize a specific map $T$, we call a connected component of $\mathcal{S}(r)$ a stable region. Note that the minimal number $n$ so that $T^{n}(I)=X(T)$ does not need to be constant within a stable region. It is clear that $[T]$ is the maximal neighbourhood $\mathcal{U}$ for which the conditions of stability hold. It is easy to prove that for $r=2$ the entire parameter space $\mathrm{ITM}(2)$ is equal to the closure of a single stability component. A very interesting open problem is to describe the geometry of stable regions (see subsection 8.2).
In the space of $\mathrm{IET}$s, the definition of stability is vacuous, as the non-wandering set is always equal to $[0,1)$, so all of the conditions for stability are trivially satisfied. On the other hand, by the Characterization of Stability Theorem 5.12, every $\mathrm{IET}$ is not stable when considered as a map in the space of $\mathrm{ITM}$s.
1.3 Summary of results
The results and pictures discussed in Subsection 1.1 suggest that the finite type maps should be prevalent in the parameter space for an arbitrary number of intervals. As mentioned at the beginning, this was already conjectured in the measure-theoretic sense by Boshernitzan and Kornfeld in the first paper on $\mathrm{ITM}$s ([MR1356616]).
Our main result shows that this conjecture is true in the topological sense, i.e. that the set of all finite type maps contains an open and dense subset of the full parameter space $\mathrm{ITM}(r)$. This is achieved by developing the theory of stability of $\mathrm{ITM}$s.
@@ 1.3 Summary of results
It is easy to show that the stable maps are of finite type (see Corollary 5.2), from which we obtain the main theorem of this paper:
@@
By Main Theorem I, the stable maps form a dense subset of $\mathrm{ITM}(r)$. Since the set $\mathcal{S}(r)$ is open by definition and stable maps are of finite type by Corollary 5.2, the result follows. ∎
@@
An eventually periodic map is the one for which every point in $I$ lands on a periodic point.
@@
For any eventually periodic map $T$, there exists an arbitrarily small perturbation $\tilde{T}$ of $T$ that is of finite type and satisfies the ACC and Matching conditions, and in particular, is stable by Theorem B.
@@
In Section 3, we prove that the eventually periodic maps are of finite type and that the set of all eventually periodic maps contains a countable dense subset of $\mathrm{ITM}(r)$. By the Characterization of Stability Theorem 5.12, we also know that the return maps to intervals of $X$ must correspond to circle rotations. The stable maps which are also eventually periodic therefore correspond to rational rotations, in the sense that the return map to any component interval of $X$ is a rational rotation. In Section 7 we show that the stable maps form a full measure subset of the set of all finite type maps.
In Subsection 8.1, we apply our results to the special family from [MR2013352] to show that the connected components of stability are exactly the coloured triangle regions in Figure 4. Moreover, any point not contained in a closure of a triangle is of infinite type. In Subsection 8.2, we will discuss the full parameter space $\mathrm{ITM}(r)$ of $\mathrm{ITM}$s on $r$ intervals, with a focus on the geometry of stable regions.
@@ 1.4 Main idea of the proof
The basic problem we deal with everywhere in the paper is: how to perturb a map $T$ so that some of its dynamical properties change, while some others do not? In analogy to smooth one-dimensional dynamics (e.g. [MR1239171]), the dynamical properties of a map depend on the behaviour of critical points. For us, the discontinuities of the map serve as the critical points. Thus the problem is about controlling critical itineraries: we want some of them to change, and some of them to stay the same up to certain return times. Because our maps are piecewise linear, this can be reformulated in terms of the linear independence of certain vectors defined by the itineraries of critical points (see Section 4 for details). The main difficulty is the following: how does the dynamics of a map $T$ prevent the linear dependence of a set of vectors? We deal with this in the proof of Theorem 4.6. This is a transversality result that is crucial in the proofs of Theorem B and Theorem C. The proof of this linear independence relies on a rather detailed and subtle dynamical analysis.
1.5 Overview of $\mathrm{IET}$s, $\mathrm{ITM}$s and piecewise isometries
The subject of interval exchange transformations ($\mathrm{IET}$s) goes back to the 60’s. Some the earliest publications were by Katok [MR0331438], Keane [MR0357739], [MR0435353] and Veech [MR0516048]. For a broad overview of $\mathrm{IET}$s, we recommend the survey papers [MR2219821], [MR2648692]. It is well-known that $\mathrm{IET}$s are connected to billiards, see for example [MR0399423] and [MR0644840], [MR0832433, Ch. 5] and [MR0889254, Ch.2, §4D], and to flows on flat surfaces, see [MR2000471], [MR644019], [MR644018], [MR1393518], [MR1733872], [MR2261104].
In Subsection 1.1 we covered the results about generic behavior and renormalization for $\mathrm{ITM}$s, so we mention the remaining topics here. The topological properties of general $\mathrm{ITM}$s were first studied by Schmeling and Troubetzkoy in [MR1796167]. Buzzi proved in [MR1855837] that all piecewise isometries on polytopes have zero topological entropy (for $\mathrm{ITM}$s this was remarked already in [MR1356616]). Buzzi and Hubert proved in [MR2054049] that all piecewise monotone maps without periodic points (with some necessary assumptions) are semi-conjugate to $\mathrm{ITM}$s (possibly with flips). They also showed that the upper bound for the number of ergodic measures of an $\mathrm{ITM}$ on $r$ intervals is $r-1$, in contrast to $\mathrm{IET}$s for which the bound is $\frac{r}{2}$ (for the precise result see [MR2648692]). This bound was realized by the family considered by Bruin in [MR2308208]. $\mathrm{ITM}$s were connected to billiards with ‘one-sided mirrors’, and analysed in [MR3449199] and [MR3403406] (see Appendix A for the explanation of this connection). Connecting $\mathrm{ITM}$s with flows on surfaces is a very interesting open problem.
$\mathrm{ITM}$s also appeared in the work of Levitt in 1993 (see [MR1231840]) in the context of group action on $\mathbb{R}$-trees. There he considered a two-parameter family of $\mathrm{ITM}$s related to the holonomy map of certain 1-dimensional foliations.